A study of the career paths of hotel general managers sent questionnaires to an SRS of 300 hotels belonging to major U.S. hotel chains. There were 183 responses. The average time these 183 general managers had spent with their current company was 10.35 years. (Take it as known that the standard deviation of time with the company for all general managers is 1.4 years.)

(a) Find the margin of error for an 80% confidence interval to estimate the mean time a general manager had spent with their current company: ______ years

(b) Find the margin of error for a 99% confidence interval to estimate the mean time a general manager had spent with their current company: ______ years

(c) In general, increasing the confidence level _________ the margin of error (width) of the confidence interval. (Enter: ''DECREASES'', ''DOES NOT CHANGE'' or ''INCREASES'', without the quotes.)
Answer #1

Ans:

a)z value for 80% confidence level=1.282

Margin of error=1.282*(1.4/sqrt(183))=0.133

b)z value for 99% confidence level=2.576

Margin of error=2.576*(1.4/sqrt(183))=0.267

c)In general,increasing the confidence level INCREASES the margin of error (width) of the confidence interval